OpenStudy (theedwardsfamily):

A bulldozer moves 640 lb of dirt in 4/5 h. Complete the unit rate. (________) lb/h

OpenStudy (theedwardsfamily):

@KamiBug @sleepyjess

OpenStudy (theedwardsfamily):

@amberxoxo

OpenStudy (theedwardsfamily):

@sammixboo

OpenStudy (sleepyjess):

640 lbs for ever 4/5 of an hour. Do you know how we would find 1/5?

OpenStudy (anonymous):

640 (ib) --------->(4/5) hour X ----------> (1) hour x=(1)*(640)/(4/5)=640/4*5=160*5= 800

OpenStudy (sleepyjess):

@AJ01 Please do not give out direct answers.

OpenStudy (theedwardsfamily):

thnx @AJ01 ! @sleepyjess thnx 2!

OpenStudy (anonymous):

direct answer would be just the answer @sleepyjess here is a method that could be used for another problems

OpenStudy (sleepyjess):

No, a direct answer is also an explanation with the answer.

OpenStudy (abhisar):

It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (theedwardsfamily):

Umm.........can I close this question now? O_0

OpenStudy (theedwardsfamily):

ooooook.....i'll close it ^_^